# Learning Model Predictive Control with Error Dynamics Regression for Autonomous Racing

### v1

@@ Abstract
This work presents a novel Learning Model Predictive Control (LMPC) strategy for autonomous racing at the handling limit that can iteratively explore and learn unknown dynamics in high-speed operational domains. We start from existing LMPC formulations and modify the system dynamics learning method. In particular, our approach uses a nominal, global, nonlinear, physics-based model with a local, linear, data-driven learning of the error dynamics. We conduct experiments in simulation, 1/10th scale hardware, and deployed the proposed LMPC on a full-scale autonomous race car used in the Indy Autonomous Challenge (IAC) with closed loop experiments at the Putnam Park Road Course in Indiana, USA. The results show that the proposed control policy exhibits improved robustness to parameter tuning and data scarcity. Incremental and safety-aware exploration toward the limit of handling and iterative learning of the vehicle dynamics in high-speed domains is observed both in simulations and experiments.
@@ I Introduction
Figure 1: Overlay of the IAC car’s trajectory running LMPC with error dynamics regression. Thw data is collected at the Putnam Park Road Course in Indiana, USA. The zoomed-in image of the hairpin turn shows iterative improvement towards optimal cornering.
In this work, we present a modification to the LMPC approach presented in [8] where we improve significantly upon the model learning strategy. In particular, we propose to learn the error dynamics between a given nominal model and the state evolution of the true system collected during run-time. This is in contrast with the prior approach which attempts to learns a model of the true system directly. We show through simulation and hardware experiments on a 1/10th scale vehicle that our approach exhibits improved robustness to parameter tuning and reduced modeling errors in scenarios when data is scarce for model learning. We finally demonstrate the effectiveness of our approach on a full-size IAC autonomous race car, the results of which are shown in Figure 1. An open source C++ implementation of our work has been made available at https://github.com/MPC-Berkeley/Racing-LMPC-ROS2.
@@
Another branch of works exploit the repetitive nature of autonomous racing in a single-vehicle scenario, and improves controller performance without any prior assumption about the distribution of the modeling error or noise. Data-driven methods, such as iterative learning control, can optimize the tracking of a repetitive trajectory given data from previous iterations by solving for control variable corrections [17]. More recently, the learning model predictive control (LMPC) framework has been proposed with a direct regression of local affine dynamics [18, 19, 8]. This combination of LMPC and regressive dynamics learning can iteratively learn the unknown system dynamics and reduce lap-time without an explicit prior reference trajectory or a nominal dynamics model. The approach has also been demonstrated on scaled vehicle hardware platforms in [18].
@@ II Problem Formulation
In this work, we model the vehicle using the planar dynamic bicycle model [20, 21] with state and input vectors:
$x$ $=[v_{x},v_{y},\omega_{z},e_{\psi},s,e_{y}]^{\top}\in\mathbb{R}^{6}$
$u$ $=[a,\delta]^{\top}\in\mathbb{R}^{2}$
where $v_{x}$, $v_{y}$, and $\omega_{z}$ are the vehicle’s longitudinal velocity, lateral velocity, and yaw rate at the center of gravity. $s$, $e_{y}$, and $e_{\psi}$ are the pose of the vehicle expressed in the Frenet frame w.r.t. a parametric path $\tau:[0,L]\mapsto\mathbb{R}^{2}$, which is assumed to be continuously twice differentiable and of total length $L$. In particular, when given an arclength $s\in[0,L]$, $\tau$ returns the global $x$-$y$ position of the path. In the case where $\tau$ forms a closed circuit, i.e. $\tau(0)=\tau(L)$, $\tau^{\prime}(0)=\tau^{\prime}(L)$, $\tau^{\prime\prime}(0)=\tau^{\prime\prime}(L)$, where $\tau^{\prime}$ and $\tau^{\prime\prime}$ are the element-wise first and second derivatives of $\tau$, we define the circuit as $\bar{\tau}(s)=\tau(\text{mod}(s,L))$ for $s\geq 0$, where mod is the modulo operator. Given a path $\tau$ and a vehicle with global position $p=(x,y)$ and heading $\psi$, the Frenet frame pose is defined as the vehicle’s path progress $s(p)=\arg\min_{s}\|\tau(s)-p\|_{2}$, its lateral deviation from the path $e_{y}(p)=\min_{s}\|\tau(s)-p\|_{2}$, and its heading deviation from the path tangent $e_{\psi}(p,\psi)=\psi-\arctan(\tau^{\prime}(s(p)))$. The inputs to the model are the longitudinal acceleration $a$ and the front steering angle $\delta$. We write the discretized nonlinear dynamics as:
@@
$\mathds{1}_{\mathcal{F}}(x)=\begin{cases}0\quad\text{if}\ x\in\mathcal{F}\\ 1\quad\text{o.w.}\end{cases}$
Due to the infinite horizon cost in (2a), it should be clear that the optimal control problem posed in (II) is not amenable to online implementation. Moreover, it is likely that there is mismatch between the vehicle model in (2c) and the true vehicle dynamics, the effects of which would be especially apparent at high speeds where the vehicle is operating at the limit of handling. To address these challenges, previous work on LMPC [19, 22] proposed a finite-horizon convex approximation of (II), which uses state and input trajectories from prior laps to synthesize a convex terminal cost-to-go function and target set. The same data is additionally used to identify affine time-varying (ATV) models which approximate the true dynamics of the vehicle. In this work, we carry over the same approach for terminal cost-to-go function and target set synthesis, but propose a modification to the system identification procedure which results in an LMPC control policy that achieves similar performance to [19, 22] w.r.t. minimum lap time, while exhibiting a significant improvement in robustness to tuning parameters.
@@ III Local Learning Model Predictive Control
The main idea of LMPC is to use the data from iterations $1$ to $j-1$ to synthesize a finite-horizon local convex approximation of (II) for iteration $j$, which is then solved in a receding horizon manner. Let us first denote the available dataset at iteration $j$ as $\mathcal{D}^{j}=(\mathbf{x}^{j},\mathbf{u}^{j})\cup\mathcal{D}^{j-1}$, where $\mathbf{x}^{j}=\{x_{0}^{j},x_{1}^{j},\dots,x_{T^{j}}^{j}\}$ and $\mathbf{u}^{j}=\{u_{0}^{j},u_{1}^{j},\dots,u_{T^{j}}^{j}\}$ are the closed-loop state and input sequence for iteration $j$ and $T^{j}$ denotes the time step at which the lap was completed for iteration $j$, i.e. the first time step where $\mathds{1}_{\mathcal{F}}(x_{k})=0$. Note that on the first iteration, the dataset $\mathcal{D}^{0}$ may be initialized using human-driven laps or closed-loop trajectories from a simple low-speed center-line tracking controller. In addition, we assume that the trajectories stored in $\mathcal{D}^{j}$ are feasible w.r.t. the constraints (2d) and reach the finish line in a finite amount of time.
@@ III-A Local Convex Target Set
As the closed-loop trajectories stored in $\mathcal{D}^{j}$ are from the actual vehicle, it is straightforward to see that if we can control the vehicle to any state in $\mathcal{D}^{j}$, then there exists a known feasible control sequence which can be used to complete the lap from that state. LMPC uses this fact to construct a safe set as a discrete collection of the states in $\mathcal{D}^{j}$, which is then used as the terminal set for a FHOCP [22]. In this work, we construct convex terminal sets which are local about a given state $x$. These terminal sets sacrifice the theoretical guarantee of persistent feasibility for the FHOCP in favor of reduced computational complexity for fast online control. In particular, for some given state $x$, we first take the $K$-nearest neighbors from the each of the previous $P$ laps by evaluating the weighted Euclidean distance over the relevant states:
@@
for some given $P$ and $D\succeq 0$. Let $\mathbf{X}^{j}(x;\mathcal{D}^{j})\in\mathbb{R}^{n\times KP}$ denote the matrix formed by the $KP$ states closest to $x$, we then define the target set as the convex hull of these states:
@@ III-B Local Terminal Cost-to-go Function
This function can be thought of as a local convex approximation to the optimal cost-to-go (from solving (II)) at iteration $j$, which when incorporated as the terminal cost-to-go function, allows an FHOCP to reason about infinite-horizon behavior of the vehicle, to the extent as it is captured in the dataset.
@@ IV Learned Vehicle Dynamics Model
We now introduce the main contribution of our work, which is a learning strategy on the dynamics states $v_{x},v_{y},\omega_{z}$ based on regressive error. The learning process take as inputs the nominal dynamics model and a set of neighboring states, and outputs a learned local affine approximation of the true vehicle dynamics.
We assume that the true dynamics of the system can be written as follows
@@
where $A^{e}$ and $B^{e}$ are the Jacobians of $e$ w.r.t. $x$ and $u$ evaluated at $\bar{z}$ and $C^{e}=e(\bar{x},\bar{u})-A^{e}\bar{x}-B^{e}\bar{u}$. It can be clearly seen from (6) that the linearization of the unknown error term is related to the system state and input $x$ and $u$, the actual state evolution $x^{+}$, and the prediction of the nominal model $\hat{x}^{+}$, which are all quantities that can be easily obtained from the dataset $\mathcal{D}^{j-1}$ at iteration $j$. As such, we query $\mathcal{D}^{j-1}$ to find the $M$ nearest neighbors $\mathbf{z}=\{z_{1},\dots,z_{M}\}$ of $\bar{z}$ and their corresponding state evolutions $\mathbf{x}^{+}=\{x_{1}^{+},\dots,x_{M}^{+}\}$ (using a technique similar to (III-A)). We then compute the prediction of nominal model at each of the state and input pairs in $\mathbf{z}$ to obtain $\hat{\mathbf{x}}^{+}=\{\hat{x}_{1}^{+},\dots,\hat{x}_{M}^{+}\}$. As mentioned before, we are interested in learning only the error dynamics associated with the velocity components of the state $x$ as the kinematic components of the state are well-understood. Therefore, using the datasets $\mathbf{z}$, $\mathbf{x}^{+}$, and $\hat{\mathbf{x}}^{+}$, we define the residuals for each velocity state as follows:
@@
$K(u)$ =$=\begin{cases}\frac{3}{4}(1-u^{2}/h^{2}),&|u|<1\\ 0,&\text{otherwise}\end{cases}.$
@@
$\Gamma_{l}^{\star}=\operatorname*{argmin}_{\Gamma_{l}}\sum_{m=1}^{M}K\left(||\bar{x}-x_{m}||^{2}_{Q}\right)y_{m}^{l}(\Gamma_{l}),$
where $Q$ denotes the relative scaling between the variables, and $y_{m}^{l}(\Gamma_{l})$ are the $l^{2}$-norms of the residuals in (7).
Finally, we can unpack the optimal regressor vectors $\Gamma_{l}^{\star}$ to construct the matrices $A^{e}$, $B^{e}$, and $C^{e}$, and add them to the nominal ATV dynamics to obtain the complete ATV model for (4c):
$A(\bar{z};\mathcal{D}^{j-1})$ $=A^{f}+\begin{bmatrix}\begin{matrix}\Gamma_{v_{x}}^{\star}[1:3]\\ \Gamma_{v_{y}}^{\star}[1:3]\\ \Gamma_{\omega_{z}}^{\star}[1:3]\\ \mathbf{0}_{3\times 3}\end{matrix}&\mathbf{0}_{6\times 3}\end{bmatrix},$
$B(\bar{z};\mathcal{D}^{j-1})$ $=B^{f}+\begin{bmatrix}\begin{matrix}\Gamma_{v_{x}}^{\star}[4]&0\\ 0&\Gamma_{v_{y}}^{\star}[4]\\ 0&\Gamma_{\omega_{z}}^{\star}[4]\end{matrix}\\ \mathbf{0}_{3\times 2}\end{bmatrix},$
$C(\bar{z};\mathcal{D}^{j-1})$ $=C^{f}+\begin{bmatrix}\Gamma_{v_{x}}^{\star}[5]&\Gamma_{v_{y}}^{\star}[5]&\Gamma_{\omega_{z}}^{\star}[5]&\mathbf{0}_{1\times 3}\end{bmatrix}^{\top},$
where $A^{f}$ and $B^{f}$ are the Jacobians of $f$ w.r.t. $x$ and $u$ evaluated at $\bar{z}$ and $C^{f}=f(\bar{x},\bar{u})-A^{f}\bar{x}-B^{f}\bar{u}$. We note that the key difference of this procedure compared to [8] is that the regressive learning is performed on the error between the prediction of the nominal model and the actual state evolution data instead of on the entire model in (5) all at once. This strategy not only provides a safety fallback when the regressive learning fails at run-time, but also demonstrates, in comparison to [8], improved robustness to parameter tuning and reduced modeling errors in scenarios when data is scarce.

### v2

@@ Abstract
This work presents a novel Learning Model Predictive Control (LMPC) strategy for autonomous racing at the handling limit that can iteratively explore and learn unknown dynamics in high-speed operational domains. We start from existing LMPC formulations and modify the system dynamics learning method. In particular, our approach uses a nominal, global, nonlinear, physics-based model with a local, linear, data-driven learning of the error dynamics. We conducted experiments in simulation and on 1/10th scale hardware, and deployed the proposed LMPC on a full-scale autonomous race car used in the Indy Autonomous Challenge (IAC) with closed loop experiments at the Putnam Park Road Course in Indiana, USA. The results show that the proposed control policy exhibits improved robustness to parameter tuning and data scarcity. Incremental and safety-aware exploration toward the limit of handling and iterative learning of the vehicle dynamics in high-speed domains is observed both in simulations and experiments.
@@ I Introduction
Figure 1: Overlay of the IAC car’s trajectory running LMPC with error dynamics regression. The data are collected at the Putnam Park Road Course in Indiana, USA. The zoomed-in image of the hairpin turn shows iterative improvement towards optimal cornering in the 1st, 5th, and 20th lap.
In this work, we present a modification to the LMPC approach presented in [8] where we improve significantly upon the model learning strategy. In particular, we propose to learn the error dynamics between a given nominal model and the state evolution of the true system collected during run-time. This is in contrast with the prior approach, which attempts to learn a model of the true system directly. We show through simulation and hardware experiments on a 1/10th scale vehicle that our approach exhibits improved robustness to parameter tuning and reduced modeling errors in scenarios when data are scarce for model learning. We finally demonstrate the effectiveness of our approach on a full-size IAC autonomous race car, the results of which are shown in Figure 1. An open source C++ implementation of our work has been made available at https://github.com/MPC-Berkeley/Racing-LMPC-ROS2.
@@
Another branch of works exploits the repetitive nature of autonomous racing in a single-vehicle scenario, and improves controller performance without any prior assumption about the distribution of the modeling error or noise. Data-driven methods, such as iterative learning control, can optimize the tracking of a repetitive trajectory given data from previous iterations by solving for control variable corrections [17]. More recently, the learning model predictive control (LMPC) framework has been proposed with a direct regression of local affine dynamics [18, 19, 8]. This combination of LMPC and regressive dynamics learning can iteratively learn the unknown system dynamics and reduce lap-time without an explicit prior reference trajectory or a nominal dynamics model. The approach has also been demonstrated on scaled vehicle hardware platforms in [18].
@@ II Problem Formulation
In this work, we model the vehicle using the planar dynamic bicycle model [20, 21] with state and input vectors $x=[v_{x},v_{y},\omega_{z},e_{\psi},s,e_{y}]^{\top}$, $u=[a,\delta]^{\top}$, where $v_{x}$, $v_{y}$, and $\omega_{z}$ are the vehicle’s longitudinal velocity, lateral velocity, and yaw rate at the center of gravity. $s$, $e_{y}$, and $e_{\psi}$ are the pose of the vehicle expressed in the Frenet frame w.r.t. a parametric path $\tau:[0,L]\mapsto\mathbb{R}^{2}$, which is assumed to be continuously twice differentiable and of total length $L$. In particular, when given an arclength $s\in[0,L]$, $\tau$ returns the global $x$-$y$ position of the path. In the case where $\tau$ forms a closed circuit, i.e., $\tau(0)=\tau(L)$, $\tau^{\prime}(0)=\tau^{\prime}(L)$, $\tau^{\prime\prime}(0)=\tau^{\prime\prime}(L)$, where $\tau^{\prime}$ and $\tau^{\prime\prime}$ are the element-wise first and second derivatives of $\tau$, we define the circuit as $\bar{\tau}(s)=\tau(\text{mod}(s,L))$ for $s\geq 0$, where mod is the modulo operator. Given a path $\tau$ and a vehicle with global position $p=(x,y)$ and heading $\psi$, the Frenet frame pose is defined as the vehicle’s path progress $s(p)=\arg\min_{s}\|\tau(s)-p\|_{2}$, its lateral deviation from the path $e_{y}(p)=\min_{s}\|\tau(s)-p\|_{2}$, and its heading deviation from the path tangent $e_{\psi}(p,\psi)=\psi-\arctan(\tau^{\prime}(s(p)))$. The inputs to the model are the longitudinal acceleration $a$ and the front steering angle $\delta$. We write the discretized nonlinear dynamics as:
@@
$\mathds{1}_{\mathcal{F}}(x)=\begin{cases}0\quad\text{if}\ x\in\mathcal{F}\\ 1\quad\text{otherwise}\end{cases}$
Due to the infinite horizon cost in (2a), it should be clear that the optimal control problem posed in (II) is not amenable to online implementation. Moreover, it is likely that there is mismatch between the vehicle model in (2c) and the true vehicle dynamics, the effects of which would be especially apparent at high speeds, where the vehicle is operating at the limit of handling. To address these challenges, previous work on LMPC [19, 22] proposed a finite-horizon convex approximation of (II), which uses state and input trajectories from prior laps to synthesize a convex terminal cost-to-go function and target set. The same data are additionally used to identify affine time-varying (ATV) models which approximate the true dynamics of the vehicle. In this work, we carry over the same approach for terminal cost-to-go function and target set synthesis, but propose a modification to the system identification procedure which results in an LMPC control policy that achieves similar performance to [19, 22] w.r.t. minimum lap time, while exhibiting a significant improvement in robustness to tuning parameters.
@@ III Local Learning Model Predictive Control
The main idea of LMPC is to use the data from iterations $1$ to $j-1$ to synthesize a finite-horizon local convex approximation of (II) for iteration $j$, which is then solved in a receding horizon manner. Let us first denote the available dataset at iteration $j$ as $\mathcal{D}^{j}=(\mathbf{x}^{j},\mathbf{u}^{j})\cup\mathcal{D}^{j-1}$, where $\mathbf{x}^{j}=\{x_{0}^{j},x_{1}^{j},\dots,x_{T^{j}}^{j}\}$ and $\mathbf{u}^{j}=\{u_{0}^{j},u_{1}^{j},\dots,u_{T^{j}}^{j}\}$ are the closed-loop state and input sequence for iteration $j$ and $T^{j}$ denotes the time step at which the lap was completed for iteration $j$, i.e., the first time step where $\mathds{1}_{\mathcal{F}}(x_{k})=0$. Note that on the first iteration, the dataset $\mathcal{D}^{0}$ may be initialized using human-driven laps or closed-loop trajectories from a simple low-speed center-line tracking controller. In addition, we assume that the trajectories stored in $\mathcal{D}^{j}$ are feasible w.r.t. the constraints (2d) and reach the finish line in a finite amount of time.
@@ III-A Local Convex Target Set
As the closed-loop trajectories stored in $\mathcal{D}^{j}$ are from the actual vehicle, it is straightforward to see that if we can control the vehicle to any state in $\mathcal{D}^{j}$, then there exists a known feasible control sequence which can be used to complete the lap from that state. LMPC uses this fact to construct a safe set as a discrete collection of the states in $\mathcal{D}^{j}$, which is then used as the terminal set for a FHOCP [22]. In this work, we construct convex terminal sets which are local about a given state $x$. These terminal sets sacrifice the theoretical guarantee of persistent feasibility for the FHOCP in favor of reduced computational complexity for fast online control. In particular, for a given $x$, we take the $K$-nearest neighbors from each of the previous $P$ laps by evaluating the weighted Euclidean distance over the relevant states:
@@ III-B Local Terminal Cost-to-go Function
This function can be thought of as a local convex approximation to the optimal cost-to-go (from solving (II)) at iteration $j$, which when incorporated as the terminal cost-to-go function, allows an FHOCP to reason about infinite-horizon behavior of the vehicle, to the extent that it is captured in the dataset.
@@ IV Learned Vehicle Dynamics Model
We now introduce the main contribution of our work, which is a learning strategy on the dynamics states $v_{x},v_{y},\omega_{z}$ based on regressive error. The learning process takes as inputs the nominal dynamics model and a set of neighboring states, and outputs a learned local affine approximation of the true vehicle dynamics. We assume that the true dynamics of the system can be written as follows
@@
where $A^{e}$ and $B^{e}$ are the Jacobians of $e$ w.r.t. $x$ and $u$ evaluated at $\bar{z}$ and $C^{e}=e(\bar{x},\bar{u})-A^{e}\bar{x}-B^{e}\bar{u}$. It can be clearly seen from (6) that the linearization of the unknown error term is related to the system state and input $x$ and $u$, the actual state evolution $x^{+}$, and the prediction of the nominal model $\hat{x}^{+}$, which are all quantities that can be easily obtained from the dataset $\mathcal{D}^{j-1}$ at iteration $j$. As such, we query $\mathcal{D}^{j-1}$ to find the $M$ nearest neighbors $\mathbf{z}=\{z_{1},\dots,z_{M}\}$ of $\bar{z}$ and their corresponding state evolutions $\mathbf{x}^{+}=\{x_{1}^{+},\dots,x_{M}^{+}\}$ (using a technique similar to (III-A)). We then compute the prediction of the nominal model at each of the state and input pairs in $\mathbf{z}$ to obtain $\hat{\mathbf{x}}^{+}=\{\hat{x}_{1}^{+},\dots,\hat{x}_{M}^{+}\}$. As mentioned before, we are interested in learning only the error dynamics associated with the velocity components of the state $x$, as the kinematic components of the state are well-understood. Therefore, using the datasets $\mathbf{z}$, $\mathbf{x}^{+}$, and $\hat{\mathbf{x}}^{+}$, we define the residuals for each velocity state as follows:
@@
$K(u)$ =$=\begin{cases}\frac{3}{4}(1-u^{2}/h^{2}),&|u|<h\\ 0,&\text{otherwise}\end{cases}.$
@@
$\Gamma_{l}^{\star}=\operatorname*{argmin}_{\Gamma_{l}}\sum_{m=1}^{M}K\left(||\bar{z}-z_{m}||^{2}_{Q}\right)y_{m}^{l}(\Gamma_{l})+\epsilon\|\Gamma_{l}\|_{2}^{2},$ (8)
where $Q\succeq 0$ denotes the relative scaling between the variables, $y_{m}^{l}(\Gamma_{l})$ are the $l^{2}$-norms of the residuals in (7), and $\epsilon>0$ is a regularization parameter.
Finally, we can construct the matrices $A^{e}$, $B^{e}$, and $C^{e}$ from $\Gamma_{l}^{\star}$ as follows:
$A^{e}$ $=\begin{bmatrix}\begin{matrix}\Gamma_{v_{x}}^{\star}[1:3]\\ \Gamma_{v_{y}}^{\star}[1:3]\\ \Gamma_{\omega_{z}}^{\star}[1:3]\\ \mathbf{0}_{3\times 3}\end{matrix}&\mathbf{0}_{6\times 3}\end{bmatrix},\ B^{e}=\begin{bmatrix}\begin{matrix}\Gamma_{v_{x}}^{\star}[4]&0\\ 0&\Gamma_{v_{y}}^{\star}[4]\\ 0&\Gamma_{\omega_{z}}^{\star}[4]\end{matrix}\\ \mathbf{0}_{3\times 2}\end{bmatrix},$
$C^{e}$ $=\begin{bmatrix}\Gamma_{v_{x}}^{\star}[5]&\Gamma_{v_{y}}^{\star}[5]&\Gamma_{\omega_{z}}^{\star}[5]&\mathbf{0}_{1\times 3}\end{bmatrix}^{\top}.$
These are then added to the nominal ATV dynamics to obtain the complete ATV model for (4c): $A(\bar{z};\mathcal{D}^{j-1})=A^{f}+A^{e}$, $B(\bar{z};\mathcal{D}^{j-1})=B^{f}+B^{e}$, and $C(\bar{z};\mathcal{D}^{j-1})=C^{f}+C^{e}$, where $A^{f}$ and $B^{f}$ are the Jacobians of $f$ w.r.t. $x$ and $u$ evaluated at $\bar{z}$ and $C^{f}=f(\bar{x},\bar{u})-A^{f}\bar{x}-B^{f}\bar{u}$. We note that the key difference between this procedure and that in [8] is that the regressive learning is performed on the error between the prediction of the nominal model and the actual state evolution data instead of on the entire model in (5).
To understand why our approach to model learning is more robust to scarce data, consider the case where all $\|\bar{z}-z_{m}\|_{Q}$ in (8) are close to or greater than the kernel bandwidth $h$ (i.e., when the data $z_{m}$ are far from the linearization point $\bar{z}$ w.r.t. $h$). When this occurs, $K(\|\bar{z}-z_{m}\|_{Q}^{2})\approx 0$, which means that the regularization term in (8) dominates the least squares objective and we can therefore conclude that the optimal solution $\Gamma_{l}^{\star}\approx 0$. When regressing the nominal dynamics matrices directly, as in [8], this would result in setting elements of those matrices to approximately zero, which would be disastrous in terms of model accuracy. On the other hand, when regressing the error matrices, setting elements to zero would simply correspond to not applying any data-based correction to the nominal linearized dynamics. This allows for a natural trade-off between the nominal and data-corrected models based on the quality of the dataset. Of course, this could be mitigated by the selection of a large enough bandwidth $h$. However, as this is a tuning parameter which is meant to act as a threshold on a weighted distance metric in high dimensions, it may not be immediately apparent what constitutes a good setting. We therefore want the LMPC policy to be robust to values of $h$ to allow for safe tuning of the controller.